A sewing teacher is buying rick-rack so her 7th grade students can trim aprons that they have made. She has 33 students and each apron requires 42 inches of trim. How many meters of rick-reck will she buy so that each student can trim one apron? Round to the nearest hundredth of a meter.
Answer by stanbon(75887)

A sewing teacher is buying rick-rack so her 7th grade students can trim aprons that they have made. She has 33 students and each apron requires 42 inches of trim. How many meters of rick-reck will she buy so that each student can trim one apron? Round to the nearest hundredth of a meter.
-----
42 inches = 0.0254*42 = 1.0668 meters
----
33*1.0668 = 35.20 meters
============================
Cheers,
Stan H.
==================